OpenStudy (anonymous):

200. mL of a 0.50 M NaOH solution requires 100.0 mL of a HCI solution to neutralize it. What is the concentration of the HCI solution? I cant remember how to do this please help me understand

OpenStudy (anonymous):

Mi*Vi=Mf*Vf (i=initial f= final) you should convers mL to L then 0.50M*0.2L=M(HCl)*0.1 M(HCL)= 1 M

OpenStudy (anonymous):

Wow that was actually simple, thank you very much for also explaining how you got the answer. Now I get it :)

OpenStudy (anonymous):

nothing :)

OpenStudy (anonymous):

Just to confirm 0.50M*0.2L = M(HCl)*0.1L 0.50M(0.2L/0.1L) = M(HCl) M(HCl) = 0.50M(2) M(HCl) = 1M